BRT now de-iced!

After the recent bad weather at the observatory we found last night that our dome had iced up. The dome could not rotate or open, forcing us to stop the system from operating last night.

Having found that the dome had iced up I enabled our dome heating system.
The gains in temperature on the green plot show the heater working to de-ice the dome.
This morning I have run a series of tests on the dome. I have found that it is operating correctly again. This means the telescope will return to operating tonight if the weather permits.

No comments:

Post a Comment